# Simple Network Mechanism Leads to Quasi-Real Brain Activation Patterns with Drosophila Connectome

### v1

@@ 4.3 Distances between Input and Activation Areas
4.4 Bilateral Response under Unilateral Stimulus
@@
In real life, the left and right hemispheres cooperate to perform complex tasks. We are interested in whether network models can simulate this cooperative effect to some extent. So we administer a unilateral stimulus to the input neurons and monitor the varying activation ratios across both hemispheres. More precisely, we independently simulate visual input neurons in the left and right hemispheres and record the rate of activated neurons in each hemisphere over time. Figure 7 illustrates the curve depicting how the response ratio evolves with each time step. It’s noteworthy that when simulating one hemisphere alone, visual processing neurons in both hemispheres are activated. At time step 5, the unstimulated hemisphere begins to have a activate pattern, which corresponds to the hypothesis of shallow connection[52]. Furthermore, our observations reveal a concurrent effect: the activation pattern in one hemisphere mirrors that in the unstimulated hemisphere.
@@
Figure 7: Experiment of activation rate under single stimulus. We set threshold $\sigma_{0}=0.8$.
Visually, the shape of the brain appears to be physically similar, but whether the network structures are similar remains unknown. To explore the principle behind similar responses caused by a unilateral stimulus, we calculated the degree, clustering coefficient, and eigenvector centrality distribution for each hemisphere. The Pearson correlation coefficient of degree, clustering coefficient, and eigenvector centrality vector is 0.9986, 0.9988, and 0.9991, respectively. According to Fig. 8, these statistical indicators for both hemispheres are extremely similar, indicating that the brain exhibits strong symmetry in network structure.
(a) Degree
(b) Clustering
(c) Eigenvector centrality
Figure 8: Degree, clustering coefficient, and eigenvector centrality distribution
5 Conclusion
Considering the insufficient evidence supporting the vitalness of network structure in brain message transmission and the effectiveness of network models in simulating brain activation patterns, we propose a large scale network communication model based on simple propagation rules. We also design an evaluation criteria to compare real and simulated activation patterns. Our research utilizes the largest adult Drosophila connectome data set, where we analyze the basic statistical properties and explore the relationship between neuron length, area, size, and connectivity degree. Experimental results demonstrate that our model produces an activation pattern across the brain that closely resembles reality when subjected to identical stimulus on input neurons. This result strongly suggests that even with simple propagation rules, network models can approximate the brain’s activation patterns. Besides, when changing the network structure, there is an incorrect activation pattern on the brain network. It is an evidence that network structure matters in generation real brain activation patterns. Furthermore, to prove network structure matters rather than propagation rules or spatial structure, we assess the network and spatial distances across different areas, discovering that the average network distance between input neurons and their corresponding activation areas is shorter. Notably, applying unilateral stimulus to input neurons elicit bilateral responses, mirroring bilateral brain coordination in reality and providing additional evidence that simple network models can closely replicate reality. Moreover, we calculate the statistic properties and find that the network structures of two hemispheres are extremely similar. We also make a large spatial network visualization software. This research reveals network models can reach the brain’s quasi-activation pattern even with simple propagation rules. Besides, it provides evidence that network structure matters in brain activity pattern generation. Other artificial neural network research may need to make their own structure close to the real network structure. However, it does not perfect the model to achieve a full-activation pattern. Future research could enable complete simulation of brain behavior by optimizing propagation rules and automatically adjusting link weights, paving the way for achieving genuine artificial intelligence.
6 Acknowledgement
This work was supported by the National Natural Science Foundation of China (72025405, 72088101), the National Social Science Foundation of China (22ZDA102), the Hunan Science and Technology Plan Project (2020TP1013, 2020JJ4673, 2023JJ40685), the Innovation Team Project of Colleges in Guangdong Province (2020KCXTD040), and the Postgraduate Scientific Research Innovation Project of Hunan Province (XJCX2023095). The authors declare that they have no conflict of interest.

### v3

@@ 2.4 Distances between Input and Activation Areas
3 Discussion
This study, through the construction of a large-scale network communication model based on the Drosophila connectome, reveals the critical role of neural network structure in generating brain-like activation patterns. Our findings suggest that models based on comprehensive neuron network data can approach the brain’s quasi-real activation patterns even with simple artificial rules. This paves the way for understanding biological intelligence, indicating that the structure of biological neural systems may offer valuable reference for designing efficient, fast-adapting intelligent systems.
Our research emphasizes the importance of network structure, compared to spatial structure or neuronal activation mechanisms, in determining brain activity patterns. This finding has significant implications for artificial neural network research, suggesting that future AI designs may need to more closely mimic real brain network structures.
Furthermore, the large spatial network visualization software we developed provides a new tool for neural network research, facilitating intuitive observation and analysis of complex neural network dynamics.
Future research directions may include optimizing propagation rules and automatically adjusting link weights to achieve complete simulation of brain behavior. These advancements will pave the way for achieving genuine artificial intelligence while deepening our understanding of the nature of biological intelligence.
4 Acknowledgement
We thank M.Murthy et al. for their invaluable help in sharing the Drosophila connectome data. This work was supported by the National Natural Science Foundation of China (72025405, 72088101), the National Social Science Foundation of China (22ZDA102), the Hunan Science and Technology Plan Project (2020TP1013, 2020JJ4673, 2023JJ40685), the Innovation Team Project of Colleges in Guangdong Province (2020KCXTD040), and the Postgraduate Scientific Research Innovation Project of Hunan Province (XJCX2023095). The authors declare that they have no conflict of interest.
APPENDIX
Appendix A Related Works
Various models have been developed to mimic or understand aspects of brain function, ranging from the cellular level to the whole brain. In this section, we review the models that aim to simulate the brain’s behavior. The review includes neural dynamic model that describes how action potentials in neurons are initiated and propagated. In addition, we introduce the large scale brain simulating methods. At last, we introduce the network communication models which refers to a computational or theoretical framework designed to understand and simulate how different parts of the brain communicate with each other.
A.1 Neuron Dynamic Model
Neuronal dynamic model is a mathematical framework designed to understand and simulate the electrical activity and dynamic behavior of neurons. These models aim to capture the fundamental mechanisms by which neurons process and transmit electrical signals through their membranes, often by using differential equations and computational techniques. In past years, neuroscientists have aimed to describing the behavior of neurons by dynamic model. One of the most famous works is the Hodgkin-Huxley (HH) model[18]. It includes a set of differential equations that describe the changes in membrane potential and the dynamics of ion channels. The key variables in the model are the membrane potential $V$, and the gating variables $m$, $n$, $h$ that represent the probability of ion channels being open.
The currents are given by:
$I_{Na}$ $=\bar{g}_{Na}m^{3}h(V-E_{Na}),$ (1)
$I_{K}$ $=\bar{g}_{K}n^{4}(V-E_{K}),$ (2)
$I_{L}$ $=\bar{g}_{L}(V-E_{L}),$ (3)
$\frac{dm}{dt}$ $=\alpha_{m}(1-m)-\beta_{m}m,$ (4)
$\frac{dn}{dt}$ $=\alpha_{n}(1-n)-\beta_{n}n,$ (5)
$\frac{dh}{dt}$ $=\alpha_{h}(1-h)-\beta_{h}h,$ (6)
Another classic theory is cable theory[19]. Cable theory is a mathematical model used to understand how electrical signals propagate along neurons, especially in their dendrites and axons. This theory applies the principles of electrical circuits to describe how the passive properties of the neuron influence signal transmission. It’s particularly useful in explaining the conduction of graded potentials, such as synaptic potentials and electrotonic (passive) spread of membrane potential changes.
$\frac{\partial^{2}V}{\partial x^{2}}=\frac{1}{r_{m}}\left(\frac{\partial V}{\partial t}+\frac{V}{r_{c}}\right),$ (7)
where $V$ is the membrane potential, $x$ is the distance along the cable, $t$ is time, $r_{m}$ is the membrane resistance per unit length, $r_{c}$ is the cytoplasmic (axial) resistance per unit length. This equation describes how the voltage $V$ changes along the length of the neuron $x$ and time $t$.
Cable theory has several important applications in neuroscience, including understanding signal decay, neural modeling and synaptic integration. Despite its usefulness, cable theory has limitations. First, real neurons have complex geometries. Cable theory don’t describe well of these complex geometries. This can lead to inaccuracies in predicting real neuronal behavior. Besides, cable theory primarily addresses passive electrical properties and does not account for active properties like action potentials, which are crucial for neural communication.
One of the simplest and most commonly used models to describe neuron dynamics in SNNs[22] is the Leaky Integrate-and-Fire (LIF) model[23]. The LIF model captures the essential characteristics of spike generation in neurons. The membrane potential $V(t)$ of a neuron at time $t$ is described by the following differential equation:
$\tau_{m}\frac{dV(t)}{dt}=-[V(t)-V_{rest}]+RI(t),$ (8)
Despite their advances, SNNs face significant challenges. One primary issue is that the SNN’s structure often draws inspiration from functional brain networks rather than accurately replicating the intricate neural connections observed at the microscale. In terms of the connection method of artificial neurons, SNNs usually use a structure similar to that of ANNs. Consequently, these models resort to architectures akin to those found in artificial neural networks, hindering their ability to authentically emulate brain functions. Additionally, SNNs are marked by their low computational efficiency, with a single forward propagation frequently requiring several seconds to complete. These limitations underscore the difficulties in considering SNNs as the optimal approach for brain simulation in their current state.
As for the large scale brain simulation model aims to simulate the whole brain’s behavior, Eugene et al.[28] presents the first large scale brain simulation model based on experimental measures in several mammalian species. The model is founded upon the global thalamocortical anatomy derived from diffusion tensor imaging of the human brain. It incorporates multiple thalamic nuclei and a six-layered cortical microcircuitry, which is grounded in three-dimensional reconstruction of single neurons from the cat visual cortex. It boasts nearly half a billion synapses, incorporating accurate receptor kinetics, short-term plasticity, and dendritic spike-timing-dependent synaptic plasticity (dendritic STDP). The model can demonstrate behavioral patterns that mimic normal brain activity. This model is a comprehensive attempt to simulate the whole neural system, which is a critical component of the brain involved in sensory processing and consciousness.
Eliasmith et al.[29] introduce a more comprehensive brain simulating model, which consists of 2,500,000 SNN neurons. This model receives only sequences of visual images and produces all its responses through a simulated arm. Despite its simplifications of real structure, the model encapsulates numerous elements of neuroanatomy, neurophysiology, and behavioral psychology, as evidenced by its performance across eight distinct tasks.
As for the standardized platform, Zeng et al. [30] presents BrainCog, a cognitive intelligence engine inspired by the brain. This platform, based on SNN, offers critical infrastructure for the creation of brain-inspired artificial intelligence and simulation of the brain. BrainCog encompasses a variety of biological neurons, encoding methods, learning principles, brain regions, and a combination of hardware and software as its core elements.
However, large scale brain simulations require the simulation of millions or even billions of neurons, posing extremely high demands on computational resources. Even the most powerful supercomputers might need weeks or even months to complete a single simulation. Additionally, as the number of neuron connections and synaptic weights increases, data storage and processing requirements also significantly rise. Besides, although large scale brain simulations aim to simulate neurons and their connections, they often fail to capture all the biological details present in the brain. Factors like neuron morphology, molecular-level dynamics, and changes in gene expression are complex biological phenomena that are difficult to fully integrate into the models. Thus, the simulation results might not fully reflect the real behavior of the brain. At last, building a large-scale brain simulation model requires a vast amount of precise biological data, such as the connectivity between neurons, synaptic strengths, types, and concentrations of neurotransmitters. However, acquiring these data is often very challenging, involving complex and expensive experimental techniques, and it’s difficult to measure comprehensively across the entire brain.
A.2 Brain Network Communication Model
The network communication model is a conceptual framework which describes how data is transmitted from one node to another over a network. It breaks down the complex process of communication into simpler, more manageable layers or components. The most widely recognized network communication models are the Open Systems Interconnection (OSI) model and the Internet Protocol Suite, commonly known as the TCP/IP model. Nowadays, scientists apply network communication models to neuron research, which are mostly used to understand how brain networks operate[33, 34, 35].
Early studies have indicated that, similar to social and electrical networks[36], brain networks also exhibit properties such as power-law distributions[37] and the rich-club phenomenon[38]. These results indicate that brain networks contain a wealth of network statistical characteristics.
Researchers designed many brain network models to describe the processes of message propagation in brain. One category of brain network communication models is the routing protocol model. The most famous model of this category is shortest path model[39], which assumes that signals between a pair of nodes travel along the shortest path between them. In the shortest path model, the shortest path length from the starting node $s$ to each node $v$, is initialized as $d(s)=0$ and $d(v)=\infty$; then, process each node in increasing order of $d(v)$, for each node, update the $d(u)$ value of its neighbors $u$. The update way can be represented as follows:
$d(u)=\min(d(u),d(v)+w(v,u)),$ (9)
where $w(v,u)$ is the weight of edge $e_{uv}$.
Information is propagated entirely through the shortest paths in these models, which means in terms of propagation cost, signal transmission is optimal. However, the key disadvantage of the model is these methods require each neuron must have knowledge about the entire network structure[40], which doesn’t correspond to the reality: every neuron only has the message of its neighbor neurons.
Another well-known model is navigation model. It suggests that neurons always send information to the neuron that can be reached with the least communication cost[41]. These models achieve the lowest communication cost. A significant achievement is that some algorithms can identify paths with 70%-100% of optimal signaling efficiency in the functional connectomes of different organisms[42]. However, the main disadvantage of navigation model is it supposes that neurons know the distance between their neighbors and a desired target region. Recently their still has a lack of evidences that reveal how this knowledge would be given to neurons in brain networks.
$P_{ij}=\frac{1}{k_{i}},$ (10)
$P_{ij}=\frac{w_{ij}}{\sum_{l\in N(i)}w_{il}},$ (11)
where $w_{ij}$ is the weight of node $i$ to node $j$,$N(i)$is the set of node $i$.
The diffusion process does not require individual neurons to have knowledge about the connectivity of connectome beyond their immediate neighbors. However, studies have shown that the efficiency of signal transmission through random walks is relatively low[44]. So it can’t describe the rapid communication process in the brain.
$a_{i}(t+1)=1\quad\text{if}\quad\frac{1}{k_{i}}\sum_{j\in N_{i}}a_{j}(t)\geq\theta_{i},$ (12)
where $N_{i}$ is the set of neighbors of node $i$, $a_{i}(t)$ is the adoption state of node $i$ at time $t$. $\theta_{i}$ is the threshold of node $i$. $k_{i}$ is the degree of node $i$. The threshold model has stronger biological interpretability because it approximates the mechanism of neuronal firing. Consequently, the threshold model has been extensively applied in artificial neural networks. However, how to set the propagation thresholds for different neurons, and how to design the dynamics model between neurons remains a challenge.
Appendix B Data and Methods
B.1 Drosophila Connectome Data Set
We employ the recently released largest Drosophila brain connectome[17], offering a thorough mapping of the entire neural network within a Drosophila female brain. This data set encompasses over 120,000 neurons and 30,000,000 synapses, including detailed information on synaptic connections between neurons. Notably, it provides high-precision 3D coordinates, accurate to the nanometer level, for both cell bodies and synapses. Additionally, this data set offers extensive labeling for various types of neurons. For a more in-depth understanding, please refer to [17][46] for detailed information. In this section, we use nodes to represent neurons and connections to represent the synapses between neurons.
Table 3 presents the basic properties of the data set. Note that we combine multiple edges between two nodes to calculate network statistic properties. The average degree before and after combining multiple edges is 501.6 and 37.01, respectively, indicating it is prevalent that the pairs of neurons are connected by multiple synapses.
Number of Nodes Number of Edges Average Degree
$131459$ $32970606$ $37.01$
$501.6$ $0.1527$ $0.0008817$
Network Diameter Assortativity Coefficient
$11$ $0.0008817$
B.2 Model
In this section, we introduce the large scale network communication model which consumes less computational resource and can efficiently utilize the network structure. To solve the problem that the link weights between neurons can’t be measured, we propose two hypotheses to estimate the link weights. In addition, we replace the node model in the network communication model with an improved LIF neuron model for comparison.
B.2.1 Symbol Systems
The original network presentation of neural systems often uses nodes to represent the neurons and using edges to represent the links (chemical synapses, electrical synapses, etc) between neurons[4]. However, this approach has a number of limitations: first, there are usually more than one link (synapse) between two neurons[49] and the links are spatially and temporally asymmetric. To illustrate, two neurons shown in Fig.S1 have multiple synapses between them, and the dynamical properties of these different synapses may vary. Additionally, due to their spatial locations, the signal transmission delay and the position of action on the downstream neurons are also different, making it an oversimplification if the synapses are merged into one edge.
Figure S1: An example of complex connection
For this reason, we use the dual method[50] on Drosophila connectome data set to model the multi-link between neurons.
In a dual graph, the nodes and edges of the original graph are interchanged, resulting in a new graph. In simple terms, a dual graph is obtained by replacing the nodes of the original graph with edges, and the edges with nodes. Fig. S2 is an example of an original graph and a dual graph. After being processed, the nodes in the dualed graph represent the synapses and links represent the cell body. Then we can use the communication model to describe the message propagation process.
(a) original graph
(b) dual graph
Figure S2: An example of original graph and dual graph
In order to standardize the symbol system, we use a edge (multi-edge, could include more than 2 nodes) to represent the neuronal cell body and use an node to represent a physical link between neurons (synapse). We regard the process of messages interactions as the nodes’ interactions via edges.
Let $N=\{n_{1},...,n_{i},...\}$ represents the set of all nodes and $E=\{e_{1},...,e_{j}...\}$ represents the sets of all edges. An obvious relationship is formalized as the following manner.
$e_{k}\cap e_{q}=\{n_{i}|(n_{i}\in e_{k})\wedge(n_{i}\in e_{q})\}.$ (13)
Note that the union of each element in $N$ and $E$ is the whole graph. The graph $G$’s definition is given as follows.
$G=\bigcup_{j}e_{j}=\bigcup_{i}\{n_{i}\}.$ (14)
Let the label of any neuron be $j$. Because neural link is directed, let ${e_{j}}^{+}$ ( ${e_{j}}^{-}$) represent the set of elements which output (input) the signals, note that ${e_{j}}$ = ${e_{j}}^{+}\cup{e_{j}}^{-}$.
Each neuron and synapse has an activate state (state = 1) or resting state (state = 0). When a neuron is at the activate state, it fires to other neurons, otherwise it keeps rest. So set the statement of the element $i$ as $X_{i}$. $X_{i}=1$ if the element $n_{i}$ is active, otherwise $X_{i}=0$. Then let $Y_{j}$ represent the statement of set $e_{j}$. $Y_{j}=1$ if the set $e_{j}$ is active, otherwise $Y_{j}=0$.
B.2.2 Large Scale Network Communication Model
By modeling the Drosophila connectome as graph model, we introduce the large scale network communication model on this graph. We use a linear threshold activation function on each neuron. Thus, general transmission function is defined in the following manner,
$Y_{j}=\sigma(\frac{1}{\beta_{j}}\sum_{i|e_{i}^{+}\cap e_{j}^{-}\neq\emptyset}\sum_{k|n_{k}\in e_{i}^{+}\cap e_{j}^{-}}\omega(n_{k},e_{i},e_{j})X_{k}),$ (15)
where $\omega(n_{k},e_{i},e_{j})$ represents the link weight between set $i$ and set $j$ (specifically, two neurons) through element $n_{k}$. Here we consider that the link weight between two neurons is not only related to the neurons themselves, but also to the synapses between them. $\sigma(.)$ is the activate function. $\beta_{j}$ is the normalization factor.
In this paper we use threshold activation function and sigmoid function, which are two .
The threshold activation function is a simple, binary decision function used in artificial neurons. It outputs a value of 1 if the input exceeds a certain threshold, and 0 otherwise. This function is also known as the Heaviside step function.
$\sigma(x)=\left\{\begin{aligned} 1\quad x\geq threshold\\ 0\quad x<threshold\\ \end{aligned}\right.$ (16)
The sigmoid activation function is a smooth, S-shaped function that maps input values to an output range between 0 and 1. It is commonly used in neural networks to introduce non-linearity and to model probability-like outputs in binary classification problems. The sigmoid function is defined as the logistic function.
$\sigma(x)=\frac{1}{1+e^{-x}}$ (17)
$X_{i}=Y_{j}\ \ \ \ if\ n_{i}\in e^{+}_{j}.$ (18)
The link weights between neurons determine the extent of influence they have on each other. Larger link weights mean a tighter ”link” between two neurons. In everyday life, the brain continuously changes the link weights between neurons in response to inputs from the external world, allowing intelligence to emerge from the micro to the macro. Therefore, if researchers want to simulate the process of information propagation in the brain on large scale in a computer, determining the link weights between neurons is essential. However, existing experimental methods are not well-suited for directly measuring the weights between neurons in a whole brain, so adopting some alternative approach is necessary. Below, we introduce a hypothesis that use different methods to estimate the connection weights between neurons.
$\omega(n_{k},e_{i},e_{j})=1.$ (19)
Additionally, We denote the sum of all link weights between neurons $i$ and $j$ as $W_{ij}$:
$W_{ij}=\sum_{k|n_{k}\in e_{i}^{+}\cap e_{j}^{-}}\omega(n_{k},e_{i},e_{j}).$ (20)
And we define $\beta_{j}$ as the maximum weight among all $W_{ij}$ related to neuron $j$. That is to say,
$\beta_{j}={{Max}(\{W_{ij}|e_{i}^{+}\cap e_{j}^{-}\neq\emptyset}\}).$ (21)
Where $Max(S)$ represents the Maximal element in set $S$.
Then the normalized equation is as follows.
$Y_{j}=\sigma(\frac{1}{\beta_{j}}\sum_{i|e_{i}^{+}\cap e_{j}^{-}\neq\emptyset}\sum_{k|n_{k}\in e_{i}^{+}\cap e_{j}^{-}}\omega(n_{k},e_{i},e_{j})X_{k}).$ (22)
B.2.3 Neuron Dynamic Model
To better reflect real-world conditions and demonstrate that the superiority of this network architecture does not solely rely on the design of the activation function, we also employed a variant of the LIF model[23] as the basic neuron model.
The membrane potential change of a neuron’s voltage can be described by the following differential equation.
$V_{j}(t+dt)=V_{j}(t)+\left(\frac{dt}{\tau}\right)\left(R\cdot I^{syn}_{j}-(V_{j}(t)-V_{\text{reset}})\right)$ (23)
Where $V_{j}$ represents the membrane potential of $j^{th}$ neuron, $R$ represents the resistance constant and $I^{syn}_{j}$ represents the total input current of neuron $j$. $\tau$ represents the time constant.
The input current $I^{\text{syn}}_{j}$ can’t be calculated directly. To estimate the value of $I^{\text{syn}}$, we make the following assumptions based on the following supposes:
1. The magnitude of the current on the synapse should be proportional to the voltage difference between the connected neurons.
2. The current on each synapse should be proportional to the connection weight between the neurons.
3. A current can only be emitted if the output neuron is in an active state.
Considering the supposes above, the equation of input current $I^{\text{syn}}_{j}$ is designed as follows:
$I^{\text{syn}}_{j}=\frac{1}{\alpha\beta_{j}}\sum_{{}_{i|e_{i}^{+}\cap e_{j}^{-}\neq\emptyset}}\sum_{k|n_{k}\in e_{i}^{+}\cap e_{j}^{-}}\left(\omega(n_{k},e_{i},e_{j})\times(V_{i}-V_{j})\times Y_{i}\right),$ (24)
where $V_{i}-V_{j}$ represents the voltage difference between neuron $i$ and $j$. $Y_{i}$ represents the state of output neuron $i$. $\beta_{j}$ is defined according to equation (21). $\alpha$ is the proportion const.
According to the LIF model, when a neuron’s voltage reaches the threshold value, the neuron will fire and then quickly resets its voltage. Based on this, the mechanism for updating the neuron voltage is designed as follows.
$\text{if}\ V_{j}\geq V_{th}&:\ Y_{j}=1\ \text{and}\ V_{j}=V_{\text{rest}}\\ \text{else}&:\ Y_{j}=0\end{aligned},$ (25)
where $V_{th}$ represents the fire voltage and $V_{\text{rest}}$ represents the rest voltage, respectively.
Other propagation mechanism is consistent with network model introduced in section B.2.2.
B.3 Evaluation Criteria
In previous study, researchers used classic indicators such as signaling cost and activation time to evaluate the quality of the model[16][51][52]. However, these indicators can’t describe the model accords with the true communication process or not, even if the activation time is short and the signaling cost is minimum, etc. To better assess whether a model can reflect the real situation of brain communication, we use the percentage of activated neurons in regions that related to (should work under) a certain stimulus. To illustrate, if the brain gets a visual stimulus, the region related to vision should work while the region not related to vision should remain a resting state. Then we record the percentage of neurons activated. If regions related to the certain function have a significant higher activation percentage than the average activation percentage of the whole brain, we call it a quasi-real activation pattern.
We define the consistency indicator $C_{r}$ in region $r$ which has a certain function (visual, olfactory, etc) is defined in the following manner.
$C_{r}=N_{r}^{act}/N_{r},$ (26)
Appendix C Introduction of Related Neurons
For further information, please read [46].
Appendix D Bilateral Response under Unilateral Stimulus
@@
In real life, the left and right hemispheres cooperate to perform complex tasks. We are interested in whether network models can simulate this cooperative effect to some extent. So we administer a unilateral stimulus to the input neurons and monitor the varying activation ratios across both hemispheres. More precisely, we independently simulate visual input neurons in the left and right hemispheres and record the rate of activated neurons in each hemisphere over time. Fig.S3 illustrates the curve depicting how the response ratio evolves with each time step. The horizontal line represents the proportion of activated neurons in all hemisphere. It is noteworthy that when simulating one hemisphere alone, visual processing neurons in both hemispheres are activated. At time step 5, the unstimulated hemisphere begins to have a activate pattern, which corresponds to the hypothesis of shallow connection[53]. Furthermore, our observations reveal a concurrent effect: the activation pattern in one hemisphere mirrors that in the unstimulated hemisphere.
@@
Figure S3: Experiment of activation rate under single stimulus. We set threshold $\sigma_{0}=0.8$.
Appendix E Impact of Perturbation Rate on Activation Areas
The main text demonstrates that with a 10% perturbation in the network, the original realistic activation patterns vanish regardless of the model used. In this section, we discuss the impact of different reconnection rates on activation patterns using the degree-preserving reconnection method.
The experiments presented in this section involve calculating patterns induced by two types of stimuli across three models under network perturbations of 1%, 3%, 5%, 10%, and 20% reconnect rates. We normalized each data point using the difference between the activation ratio of region $r$ and the average activation ratio $\bar{C}$ of the entire brain, that is to say, $\Delta C=C_{r}-\bar{C}$.
The experimental results are shown in Fig.S4, where 0 represents the original, unperturbed experimental results. It can be observed that at the 0 point, there is a significant difference in the relative activation ratios between visual and olfactory areas under specific stimulus. However, this significant realistic activation patterns which we discussed in the main text disappear with a 1% network perturbation, replaced by a roughly equal activation ratio across all regions. This indicates that even a 1% change in network structure can lead to a rapid loss of network functionality specificity. Additionally, it is noted that small-scale network perturbations (1%-5%) have a lesser impact on olfactory stimulus compared to visual stimulus. This is because the number of neurons associated with olfactory functions is significantly fewer than those related to visual functions, making it less likely for small-scale network structure changes to affect these neurons, thereby leading to a smaller impact on functionality.
At the same time, various horizontal comparisons indicate that while there are no significant differences among different models in forming activation patterns, more complex neuron models have a stronger ability to maintain their original activation patterns under network disturbances. This is because the signals generated by complex neuron models have greater redundancy and functional diversity. For instance, the LIF (Leaky Integrate-and-Fire) model has a voltage decay mechanism, where the voltage gradually decays in the absence of a signal. Even if some input signals are erroneous, as long as the proportion of errors is not very large, the decay mechanism can filter out the impact of these erroneous signals.
These results indicate that the formation of realistic activation patterns is highly sensitive to changes in network structure, with even minor perturbations capable of disrupting this network configuration. This indirectly substantiates that the authentic network structure is crucial for the formation of genuine activation patterns. Additionally, it highlights that complex functions (such as vision) are more susceptible to network disturbances.
(a) threshold model, visual stimulus
(b) sigmoid model, visual stimulus
(c) LIF model, visual stimulus
(d) threshold model, olfactory stimulus
(e) sigmoid model, olfactory stimulus
(f) LIF model, olfactory stimulus
Figure S4: Relationships of activation rate with reconnect rate. We conduct 5 repeated experiments and take the average and standard deviation.
Appendix F Large Spatial Network Visualization
The visualization software of experiment observation on large spatial networks is important for researchers[57][58]. However, there is still a lack of visualization system that can observe information propagation on large spatial networks in real-time, which poses difficulties in experiment. To observe directly the communication progress on this super large Drosophila connectome, we develop a large spatial network visualization software. This software will be made available to other researchers with the author’s permission.
The drawing part mainly employs the Canvas provided by HTML5 and we use the Three.js based on WebGL technology to achieve spatial network visualization effects, while the user interface display is constructed by Vue.js, providing users with an intuitive and interactive interface to display and analyze data.
F.1 Front End Technical Details
F.2 Back End Technical Details
F.3 Function Display
Fig.SS5 is an example of the 3D visualization system.
(a) Side view 1
(b) Side view 2
(c) Side view 3
(d) Top view 1
(e) Top view 2
(f) Top view 3
(g) Front view 1
(h) Front view 2
(i) Front view 3
Figure S5: Visualization of Spatial Drosophila Brain Network
All nodes’ colors can be adjusted by visible selection or input RGB code (Fig.SS6).
Figure S6: Display of color selection
Besides, from square to round, the shape of nodes can be adjusted even real neuron (Fig.SS7).
Figure S7: Display of color selection
It also can present different single areas (Fig.SS8).
Figure S8: Display of color selection